OpenStudy (anonymous):

If a 픠 b = a -b( a - 2b), then what is x 픠 3

OpenStudy (anonymous):

the is a Smybolism Equation

OpenStudy (anonymous):

this not the

myininaya (myininaya):

\[x-3(x-2(3))=x-3(x-6)\]

myininaya (myininaya):

x-3x+18 -2x+18

OpenStudy (anonymous):

oh right forgot the distributive property

OpenStudy (anonymous):

I'll have to go look that up, never heard of it...

OpenStudy (anonymous):

O_O Never heard of the distributive property ?!?!

OpenStudy (anonymous):

No not that, a symbolism equation.

OpenStudy (anonymous):

I assume u mean symbolic equation, still don't know what that funny u thingy is...

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

what funny thing? who's laughing?

OpenStudy (anonymous):

that funny u...

OpenStudy (anonymous):

what....?

OpenStudy (anonymous):

oh that

OpenStudy (anonymous):

i tried to enter a box looking math symbol but that's what came up in openstudy

myininaya (myininaya):

well that symbol is telling us what operation to perform

OpenStudy (anonymous):

you mean it's an unknown operator, that's all?

OpenStudy (anonymous):

Just sub x for a and 3 for b?

myininaya (myininaya):

the operator lets call it o a o b= a-b(a-2b) o is an operation that is being defined in terms of operations that we do know

myininaya (myininaya):

x o 3 =x-3(x-2(3))

OpenStudy (anonymous):

k, clear. Tks.

myininaya (myininaya):

x o 3=x-3x+18=-2x+18 3 o x=3-x(3-2x)=3-3x+2x^2 so obviously the operation is not commutative x o 3 does not equal 3 o x